a baseball team has home games on Thursday and Saturday. the two games together earn 4791.00for the team. Thursdays game generates 269.00 less then Saturdays games.How much money was taken in at each game?
Answer by checkley77(12844)

X+(X-269)=4791
2X=4791+269
2X=5060
X=506C/2
X=$2530 FOR THE SATURDAY GAME
2530-269=$2261 FOR THE THURSDAY GAME.
PROOF:
2530+2261=4791
4781=4791